OpenStudy (anonymous):

Complete the table of values for the following equation. 11x + 3y = -57 x-values y-values -3 0 -19 3 6 9 -52 Help Me Please ! i'll fan & give you points

OpenStudy (anonymous):

you need to plug in the values for "x" and solve for "y". i.e, 3y = 11x-57 For x=-3, 3y = 11*(-3)-57 = -33-57 = -90 or y = -90/3 = -30. You can find the value of y similarly for other values of x.